SHOW YOUR WORK

The human body is composed of water, proteins, fat, and minerals. Its

specific heat reflects this composition. With 72% water and 28% protein, calculate the specific heat of the human body.

Substance Specific heat(cal/g°C) 1 Water Ice 0.480 Soil 0.2 to 0.8, depending on water content 0.214 0.4 Aluminum Protein
Answer #1

So net specific heat of human body is

C_p=\frac{C_w(\% of \ \ water)+C_{pr}(\% of \ \ protein)}{100}

so we get

C_p=\frac{72(1)+28(0.4)}{100}=0.832 Cal/\degree Cg